OpenStudy (anonymous):

Use the Associative Property of Addition to find the total of 1, 10, and 18 in two different ways.

hartnn (hartnn):

(1+10)+18 = 11+18 = 29 1+ (10+18) = 1+28 =29

OpenStudy (anonymous):

Basically, you have to rearrange the numbers any two ways.